# GDSR: Global-Detail Integration through Dual-Branch Network with Wavelet Losses for Remote Sensing Image Super-Resolution

### v1

@@ IV-C1 Effect of Key Modules
1 | 28.89/0.7633 | 0.3565 | 3.65M | 211.64G | 24.6 |
2 | 28.97/0.7663 | 0.3514 | 6.82M | 395.14G | 12.6 |
4 | 29.03/0.7686 | 0.3476 | 13.17M | 762.14G | 6.4 |
6 | 29.03/0.7686 | 0.3463 | 19.52M | 1129.14G | 4.3 |
8 | 29.04/0.7692 | 0.3456 | 25.86M | 1496.14G | 3.2 |
10 | 29.05/0.7695 | 0.3450 | 32.21M | 1863.13G | 2.5 |
Methods | Venue | AID [39] | AID_CDM [39] | RSSRD-QH | RSSRD-QH_CDM | ||||||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||
Bicubic | - | 27.89 | 0.7238 | 0.4400 | 23.15 | 0.5238 | 0.6911 | 26.45 | 0.6447 | 0.4243 | 22.34 | 0.3525 | 0.6993 |
SRCNN [6] | CVPR’2014 | 28.22 | 0.7363 | 0.3973 | 23.38 | 0.5455 | 0.6569 | 26.54 | 0.6512 | 0.4241 | 22.44 | 0.3588 | 0.7523 |
SRGAN [8] | CVPR’2017 | 28.80 | 0.7597 | 0.3601 | 23.78 | 0.5648 | 0.6314 | 26.68 | 0.6586 | 0.4170 | 22.64 | 0.3759 | 0.7421 |
EDSR [9] | CVPR’2017 | 28.94 | 0.7653 | 0.3513 | 23.91 | 0.5698 | 0.6139 | 26.70 | 0.6598 | 0.4141 | 22.74 | 0.3817 | 0.7283 |
SwinIR [14] | CVPR’2021 | 28.95 | 0.7658 | 0.3512 | 23.89 | 0.5705 | 0.6169 | 26.70 | 0.6597 | 0.4133 | 22.72 | 0.3780 | 0.7432 |
HAT [15] | CVPR’2023 | 28.99 | 0.7671 | 0.3479 | 23.99 | 0.5737 | 0.6080 | 26.71 | 0.6602 | 0.4111 | 22.79 | 0.3823 | 0.7312 |
MambaIR [18] | ECCV’2024 | 29.02 | 0.7684 | 0.3487 | 24.07 | 0.5763 | 0.6000 | 26.72 | 0.6607 | 0.4124 | 22.84 | 0.3835 | 0.7387 |
ConvFormerSR [16] | TGRS’2024 | 28.90 | 0.7637 | 0.3516 | 23.97 | 0.5710 | 0.6118 | 26.69 | 0.6603 | 0.4090 | 22.77 | 0.3779 | 0.7365 |
FreMamba [20] | TMM’2024 | 28.96 | 0.7672 | 0.3410 | 23.89 | 0.5692 | 0.6148 | 26.70 | 0.6601 | 0.4112 | 22.60 | 0.3660 | 0.7242 |
GDSR_TC (Ours) | - | 28.96 | 0.7660 | 0.3500 | 24.01 | 0.5717 | 0.6126 | 26.71 | 0.6604 | 0.4144 | 22.79 | 0.3776 | 0.7401 |
GDSR_MC (Ours) | - | 28.97 | 0.7668 | 0.3472 | 24.03 | 0.5734 | 0.6080 | 26.71 | 0.6602 | 0.4107 | 22.84 | 0.3819 | 0.7327 |
GDSR (Ours) | - | 29.03 | 0.7686 | 0.3476 | 24.10 | 0.5770 | 0.5986 | 26.72 | 0.6612 | 0.4106 | 22.86 | 0.3861 | 0.7268 |
GDSR_L (Ours) | - | 29.04 | 0.7692 | 0.3456 | 24.15 | 0.5795 | 0.5936 | 26.73 | 0.6615 | 0.4098 | 22.89 | 0.3875 | 0.7253 |
Methods | Venue | AID [39] | AID_CDM [39] | RSSRD-QH | RSSRD-QH_CDM | ||||||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||
SRCNN [6] | CVPR’2014 | 28.25 | 0.7369 | 0.3998 | 23.38 | 0.5456 | 0.6609 | 26.55 | 0.6498 | 0.4221 | 22.44 | 0.3590 | 0.7485 |
SRGAN [8] | CVPR’2017 | 28.82 | 0.7605 | 0.3672 | 23.80 | 0.5652 | 0.6458 | 26.69 | 0.6581 | 0.4183 | 22.65 | 0.3760 | 0.7456 |
EDSR [9] | CVPR’2017 | 28.96 | 0.7658 | 0.3524 | 23.94 | 0.5711 | 0.6233 | 26.70 | 0.6596 | 0.4128 | 22.76 | 0.3821 | 0.7357 |
SwinIR [14] | CVPR’2021 | 28.98 | 0.7668 | 0.3533 | 23.90 | 0.5704 | 0.6301 | 26.71 | 0.6592 | 0.4147 | 22.73 | 0.3781 | 0.7488 |
HAT [15] | CVPR’2023 | 29.04 | 0.7692 | 0.3483 | 24.01 | 0.5736 | 0.6183 | 26.72 | 0.6597 | 0.4137 | 22.81 | 0.3829 | 0.7348 |
MambaIR [18] | ECCV’2024 | 29.04 | 0.7690 | 0.3521 | 24.08 | 0.5767 | 0.6114 | 26.72 | 0.6598 | 0.4134 | 22.85 | 0.3842 | 0.7388 |
ConvFormerSR [16] | TGRS’2024 | 28.93 | 0.7647 | 0.3524 | 23.99 | 0.5707 | 0.6281 | 26.71 | 0.6591 | 0.4141 | 22.77 | 0.3770 | 0.7427 |
FreMamba [20] | TMM’2024 | 28.94 | 0.7673 | 0.3397 | 23.89 | 0.5682 | 0.6318 | 26.70 | 0.6599 | 0.4130 | 22.64 | 0.3698 | 0.7130 |
GDSR_TC (Ours) | - | 28.99 | 0.7670 | 0.3522 | 24.02 | 0.5719 | 0.6237 | 26.72 | 0.6595 | 0.4162 | 22.79 | 0.3776 | 0.7437 |
GDSR_MC (Ours) | - | 29.00 | 0.7675 | 0.3512 | 24.06 | 0.5742 | 0.6159 | 26.72 | 0.6591 | 0.4122 | 22.84 | 0.3815 | 0.7395 |
GDSR (Ours) | - | 29.04 | 0.7689 | 0.3522 | 24.12 | 0.5774 | 0.6121 | 26.73 | 0.6603 | 0.4116 | 22.88 | 0.3862 | 0.7313 |
GDSR_L (Ours) | - | 29.06 | 0.7694 | 0.3499 | 24.16 | 0.5791 | 0.6069 | 26.73 | 0.6603 | 0.4119 | 22.90 | 0.3875 | 0.7314 |
Fig. 8 provided visual comparisons under different branches and reconstruction strategies. It was evident that the RGEG branch had a significantly larger receptive field than the RDEG branch and activated more widely distributed pixels for image reconstruction. As a result, the RGEG branch achieved better PSNR and SSIM values. Two Global-Detail reconstruction strategies, F2B and F2D, yielded different results. F2B, which sent the reconstructed features back to both RGEG and RDEG branches, largely retained the wide receptive field benefits of the RGEG branch and similar DI values. However, its SR results were comparable to those of the single RGEG branch. In contrast, F2D, which sent the reconstructed features only back to the RDEG branch, achieved DI values between those of the RGEG and RDEG branches. The SR results of the F2D strategy showed that it correctly reconstructed textures and delivered more desirable outcomes. This finding aligned with previous studies, which suggested that while expanding the model’s receptive field allowed consideration of a broader range of pixels, it could also lead to incorrect texture reconstruction. This indicated that simply leveraging information from more pixels was insufficient for accurate reconstruction; effectively utilizing pixel information within the receptive field was crucial. As shown in the LAM Attribution map in Fig. 8, the F2D strategy, compared to F2B and the RGEG branch, demonstrated a greater contribution of pixels within the corresponding regions to the SR results. In contrast, the distant pixels captured by the wide receptive field of the RGEG branch (e.g., pixels in the upper part of the figure) contributed less to the SR results. This suggested that F2D, by routing features exclusively back to the RDEG branch, enabled the model to utilize information more efficiently and reconstruct accurate textures by correctly leveraging the pixels within the receptive field.
(d) Effect of Network Depth on Performance. To investigate the effect of network depth on the performance of GDSR, we varied the shared depth of both the RDEG and RGEG branches while keeping other design parameters constant. As illustrated in Table IV, the experimental outcomes consistently demonstrated an enhancement in performance with deeper networks. Moreover, an evaluation of the trade-offs in terms of computational complexity was conducted. As the network depth increased, the model’s parameter count, FLOPS, and FPS were also scaled accordingly. These results supported the effectiveness of GDSR’s dual-branch design, where the shared depth allowed for both high-quality super-resolution and manageable computational overhead.

### v4

@@ IV-C1 Effect of Key Modules
| 1 | 28.89/0.7633 | 0.3565 | 3.65M | 94.06G | 53.0 |
| 2 | 28.97/0.7663 | 0.3514 | 6.82M | 175.62G | 27.3 |
| 4 | 29.03/0.7686 | 0.3476 | 13.17M | 338.73G | 14.8 |
| 6 | 29.03/0.7686 | 0.3463 | 19.52M | 501.85G | 9.3 |
| 8 | 29.04/0.7692 | 0.3456 | 25.86M | 664.96G | 7.0 |
| 10 | 29.05/0.7695 | 0.3450 | 32.21M | 828.07G | 5.6 |
IV-C2 Effect of Wavelet Loss
Ablation experiments were conducted across multiple configurations to evaluate the effectiveness of incorporating wavelet loss in training. Quantitative results (Table VI) demonstrate that our full configuration achieves the highest PSNR and SSIM while attaining the lowest LPIPS, outperforming both baselines and the dual-group-only variant. Visual evidence in Fig. 9 further corroborates these findings, the inclusion of the proposed Dual-Group Multi-Scale Wavelet Loss significantly improved the reconstruction, particularly in the high-frequency components. This improvement is attributed to the synergistic integration of dual-coefficient grouping strategy and cross-resolution frequency consistency, which collectively overcome the rigidity of fixed-scale wavelet methods that process subbands in isolation. These results highlight Wavelet Loss’s role in guiding the model to recover intricate structures and accurately reconstruct complex textures, thereby enhancing overall image quality.
However, as evidenced in Table VI, all wavelet loss variants exhibit higher LPIPS values compared to the baseline, indicating an inherent challenge in reconciling frequency-domain fidelity with perceptual authenticity. Specifically, the uniform coefficient configuration [41] yields the worst LPIPS due to its undifferentiated weighting. Our dual-group multi-scale design strategically mitigates this tradeoff. While the proposed method’s LPIPS remains marginally higher than , it achieves the lowest LPIPS among wavelet-based approaches. The resultant configuration delivers superior PSNR/SSIM with minimal perceptual compromise, demonstrating that while frequency-domain constraints inherently elevate LPIPS, our method optimally balances this tradeoff with substantial PSNR/SSIM gains.
| Method | PSNR (dB) | SSIM | LPIPS |
| loss | 29.0329 | 0.7686 | 0.3476 |
| Baseline [41]: Uniform coefficients | 29.0182 | 0.7676 | 0.3565 |
| Baseline [41]: Asymmetric coefficients | 29.0413 | 0.7687 | 0.3534 |
| Proposed: Dual-Group | 29.0468 | 0.7690 | 0.3535 |
| Proposed: Dual-Group Multi-Scale | 29.0514 | 0.7694 | 0.3519 |
| Methods | Type | AID x2 | AID x3 | AID x4 | |||||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
| Bicubic |
30.48 | 0.8269 | 0.2769 | 27.89 | 0.7238 | 0.4400 | 26.43 | 0.6541 | 0.5640 | |||
| SRCNN [6] | 30.87 | 0.8352 | 0.2711 | 28.22 | 0.7363 | 0.3973 | 26.68 | 0.6679 | 0.4862 | ||||
| EDSR [9] | 31.70 | 0.8560 | 0.2416 | 28.94 | 0.7653 | 0.3513 | 27.30 | 0.6994 | 0.4381 | ||||
| SRGAN [8] | 31.57 | 0.8523 | 0.2481 | 28.80 | 0.7597 | 0.3601 | 27.20 | 0.6939 | 0.4466 | ||||
| SwinIR [12] | 31.71 | 0.8567 | 0.2370 | 28.95 | 0.7658 | 0.3512 | 27.32 | 0.6993 | 0.4373 | ||||
| HAT [13] | 31.77 | 0.8579 | 0.2351 | 28.99 | 0.7671 | 0.3479 | 27.34 | 0.7004 | 0.4342 | ||||
| HAUNet [25] | 31.48 | 0.8507 | 0.2503 | 28.75 | 0.7580 | 0.3657 | 27.15 | 0.6913 | 0.4523 | ||||
| SPT [27] | 31.61 | 0.8547 | 0.2391 | 28.86 | 0.7620 | 0.3553 | 27.26 | 0.6962 | 0.4418 | ||||
| ConvFormerSR [28] | 31.64 | 0.8551 | 0.2298 | 28.90 | 0.7637 | 0.3516 | 27.26 | 0.6964 | 0.4396 | ||||
| MambaIR [15] | 31.78 | 0.8583 | 0.2354 | 29.02 | 0.7684 | 0.3487 | 27.37 | 0.7019 | 0.4348 | ||||
| FreMamba [16] | 31.62 | 0.8558 | 0.2215 | 28.96 | 0.7672 | 0.3410 | 27.32 | 0.7010 | 0.4255 | ||||
| GDSR (Ours) | 31.81 | 0.8592 | 0.2366 | 29.05 | 0.7694 | 0.3519 | 27.39 | 0.7022 | 0.4376 | ||||
| Bicubic |
24.11 | 0.5545 | 0.6046 | 23.15 | 0.5238 | 0.6911 | 22.54 | 0.5091 | 0.7543 | |||
| SRCNN [6] | 24.44 | 0.5854 | 0.5836 | 23.38 | 0.5455 | 0.6569 | 22.70 | 0.5245 | 0.7100 | ||||
| EDSR [9] | 24.90 | 0.6153 | 0.5341 | 23.91 | 0.5698 | 0.6139 | 23.16 | 0.5444 | 0.6753 | ||||
| SRGAN [8] | 24.78 | 0.6070 | 0.5609 | 23.78 | 0.5648 | 0.6314 | 23.02 | 0.5396 | 0.6909 | ||||
| SwinIR [12] | 24.95 | 0.6144 | 0.5353 | 23.89 | 0.5705 | 0.6169 | 23.17 | 0.5452 | 0.6737 | ||||
| HAT [13] | 25.06 | 0.6183 | 0.5276 | 23.99 | 0.5737 | 0.6080 | 23.26 | 0.5475 | 0.6671 | ||||
| HAUNet [25] | 25.03 | 0.6169 | 0.5287 | 23.90 | 0.5692 | 0.6185 | 23.19 | 0.5441 | 0.6783 | ||||
| SPT [27] | 24.86 | 0.6086 | 0.5457 | 23.88 | 0.5679 | 0.6257 | 23.20 | 0.5447 | 0.6749 | ||||
| ConvFormerSR [28] | 25.04 | 0.6137 | 0.5378 | 23.97 | 0.5710 | 0.6118 | 23.23 | 0.5448 | 0.6742 | ||||
| MambaIR [15] | 25.17 | 0.6230 | 0.5120 | 24.07 | 0.5763 | 0.6000 | 23.32 | 0.5497 | 0.6590 | ||||
| FreMamba [16] | 25.00 | 0.6227 | 0.4977 | 23.89 | 0.5692 | 0.6148 | 23.06 | 0.5403 | 0.6735 | ||||
| GDSR (Ours) | 25.28 | 0.6256 | 0.5225 | 24.12 | 0.5771 | 0.6106 | 23.39 | 0.5510 | 0.6664 | ||||
| UCMerced x2 | UCMerced x3 | UCMerced x4 | |||||||||||
| Bicubic |
31.10 | 0.8892 | 0.2013 | 27.59 | 0.7731 | 0.3548 | 25.73 | 0.6813 | 0.4838 | |||
| SRCNN [6] | 32.33 | 0.9102 | 0.1635 | 28.48 | 0.8028 | 0.2919 | 26.39 | 0.7137 | 0.3944 | ||||
| EDSR [9] | 34.31 | 0.9337 | 0.0951 | 29.98 | 0.8464 | 0.2048 | 27.70 | 0.7644 | 0.3002 | ||||
| SRGAN [8] | 34.07 | 0.9317 | 0.1007 | 29.79 | 0.8427 | 0.2207 | 27.53 | 0.7571 | 0.3181 | ||||
| SwinIR [12] | 34.35 | 0.9344 | 0.0942 | 30.06 | 0.8514 | 0.2018 | 27.75 | 0.7684 | 0.2956 | ||||
| HAT [13] | 34.30 | 0.9339 | 0.0936 | 30.08 | 0.8527 | 0.1979 | 27.77 | 0.7700 | 0.2914 | ||||
| HAUNet [25] | 33.68 | 0.9270 | 0.1113 | 29.47 | 0.8343 | 0.2333 | 27.25 | 0.7479 | 0.3381 | ||||
| SPT [27] | 34.12 | 0.9319 | 0.0968 | 29.79 | 0.8439 | 0.2115 | 27.55 | 0.7613 | 0.3072 | ||||
| ConvFormerSR [28] | 34.27 | 0.9330 | 0.0945 | 29.89 | 0.8461 | 0.2123 | 27.59 | 0.7614 | 0.3109 | ||||
| MambaIR [15] | 34.52 | 0.9356 | 0.0898 | 30.14 | 0.8513 | 0.1934 | 27.75 | 0.7694 | 0.2874 | ||||
| FreMamba [16] | 33.87 | 0.9283 | 0.1053 | 29.95 | 0.8471 | 0.2025 | 27.61 | 0.7639 | 0.2992 | ||||
| GDSR (Ours) | 34.53 | 0.9354 | 0.0918 | 30.15 | 0.8500 | 0.2057 | 27.79 | 0.7683 | 0.3099 | ||||
| Bicubic |
22.46 | 0.5128 | 0.6323 | 21.35 | 0.4678 | 0.7330 | 20.86 | 0.4558 | 0.7846 | |||
| SRCNN [6] | 22.71 | 0.5392 | 0.6277 | 21.48 | 0.4860 | 0.7077 | 20.96 | 0.4702 | 0.7593 | ||||
| EDSR [9] | 22.99 | 0.5635 | 0.5485 | 21.67 | 0.5056 | 0.6402 | 21.10 | 0.4850 | 0.6917 | ||||
| SRGAN [8] | 22.96 | 0.5603 | 0.5675 | 21.65 | 0.5041 | 0.6569 | 21.10 | 0.4836 | 0.7051 | ||||
| SwinIR [12] | 23.09 | 0.5682 | 0.5496 | 21.77 | 0.5096 | 0.6481 | 21.18 | 0.4882 | 0.6900 | ||||
| HAT [13] | 23.22 | 0.5758 | 0.5338 | 21.82 | 0.5129 | 0.6374 | 21.18 | 0.4895 | 0.6851 | ||||
| HAUNet [25] | 23.03 | 0.5620 | 0.5635 | 21.69 | 0.5033 | 0.6655 | 21.07 | 0.4824 | 0.7117 | ||||
| SPT [27] | 23.22 | 0.5724 | 0.5464 | 21.82 | 0.5083 | 0.6565 | 21.26 | 0.4884 | 0.7055 | ||||
| ConvFormerSR [28] | 23.30 | 0.5739 | 0.5453 | 21.91 | 0.5124 | 0.6459 | 21.27 | 0.4898 | 0.6951 | ||||
| MambaIR [15] | 23.23 | 0.5778 | 0.5254 | 21.76 | 0.5129 | 0.6247 | 21.16 | 0.4903 | 0.6749 | ||||
| FreMamba [16] | 23.41 | 0.5783 | 0.5287 | 21.73 | 0.5129 | 0.6200 | 21.12 | 0.4894 | 0.6723 | ||||
| GDSR (Ours) | 23.45 | 0.5836 | 0.5304 | 21.93 | 0.5156 | 0.6323 | 21.23 | 0.4894 | 0.6828 | ||||
| Methods | Type | RSSRD-QH x2 | RSSRD-QH x3 | RSSRD-QH x4 | |||||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
| Bicubic |
29.37 | 0.8149 | 0.2564 | 26.45 | 0.6447 | 0.4243 | 24.92 | 0.5224 | 0.5764 | |||
| SRCNN [6] | 29.58 | 0.8218 | 0.3028 | 26.54 | 0.6512 | 0.4241 | 24.98 | 0.5268 | 0.5613 | ||||
| EDSR [9] | 29.83 | 0.8299 | 0.2838 | 26.70 | 0.6598 | 0.4141 | 25.10 | 0.5343 | 0.5487 | ||||
| SRGAN [8] | 29.80 | 0.8293 | 0.2859 | 26.68 | 0.6586 | 0.4170 | 25.08 | 0.5333 | 0.5522 | ||||
| SwinIR [12] | 29.83 | 0.8305 | 0.2844 | 26.70 | 0.6597 | 0.4133 | 25.10 | 0.5349 | 0.5474 | ||||
| HAT [13] | 29.84 | 0.8307 | 0.2836 | 26.71 | 0.6602 | 0.4111 | 25.10 | 0.5352 | 0.5454 | ||||
| HAUNet [25] | 29.79 | 0.8289 | 0.2888 | 26.65 | 0.6569 | 0.4177 | 25.07 | 0.5326 | 0.5494 | ||||
| SPT [27] | 29.83 | 0.8306 | 0.2875 | 26.69 | 0.6602 | 0.4161 | 25.10 | 0.5349 | 0.5478 | ||||
| ConvFormerSR [28] | 29.82 | 0.8310 | 0.2786 | 26.69 | 0.6603 | 0.4090 | 25.10 | 0.5354 | 0.5452 | ||||
| MambaIR [15] | 29.86 | 0.8311 | 0.2825 | 26.72 | 0.6607 | 0.4124 | 25.11 | 0.5359 | 0.5445 | ||||
| FreMamba [16] | 29.80 | 0.8300 | 0.2811 | 26.70 | 0.6601 | 0.4112 | 25.11 | 0.5366 | 0.5414 | ||||
| GDSR (Ours) | 29.87 | 0.8310 | 0.2823 | 26.73 | 0.6601 | 0.4130 | 25.12 | 0.5354 | 0.5465 | ||||
| Bicubic |
23.10 | 0.3991 | 0.5996 | 22.34 | 0.3525 | 0.6993 | 21.85 | 0.3284 | 0.7764 | |||
| SRCNN [6] | 23.27 | 0.4159 | 0.6520 | 22.44 | 0.3588 | 0.7523 | 21.91 | 0.3311 | 0.8078 | ||||
| EDSR [9] | 23.67 | 0.4485 | 0.6269 | 22.74 | 0.3817 | 0.7283 | 22.15 | 0.3444 | 0.8028 | ||||
| SRGAN [8] | 23.57 | 0.4387 | 0.6442 | 22.64 | 0.3759 | 0.7421 | 22.09 | 0.3408 | 0.8174 | ||||
| SwinIR [12] | 23.68 | 0.4463 | 0.6232 | 22.72 | 0.3780 | 0.7432 | 22.15 | 0.3440 | 0.8104 | ||||
| HAT [13] | 23.75 | 0.4528 | 0.6120 | 22.79 | 0.3823 | 0.7312 | 22.18 | 0.3456 | 0.8061 | ||||
| HAUNet [25] | 23.72 | 0.4501 | 0.6136 | 22.76 | 0.3797 | 0.7360 | 22.16 | 0.3442 | 0.8096 | ||||
| SPT [27] | 23.58 | 0.4353 | 0.6512 | 22.73 | 0.3743 | 0.7520 | 22.14 | 0.3415 | 0.8144 | ||||
| ConvFormerSR [28] | 23.67 | 0.4398 | 0.6267 | 22.77 | 0.3779 | 0.7365 | 22.18 | 0.3436 | 0.8051 | ||||
| MambaIR [15] | 23.82 | 0.4568 | 0.6036 | 22.84 | 0.3835 | 0.7387 | 22.23 | 0.3476 | 0.8003 | ||||
| FreMamba [16] | 23.59 | 0.4329 | 0.6009 | 22.60 | 0.3660 | 0.7242 | 22.06 | 0.3358 | 0.7894 | ||||
| GDSR (Ours) | 23.75 | 0.4454 | 0.6197 | 22.87 | 0.3864 | 0.7308 | 22.27 | 0.3488 | 0.8044 | ||||